# Towards Robust Automated Perceptual Voice Quality Assessment with Speech Foundation Models

### v2

@@ Abstract
Perceptual voice quality assessment is essential for diagnosing and monitoring voice disorders by providing standardized evaluations of vocal function. Traditionally, expert raters use standard scales such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and Grade, Roughness, Breathiness, Asthenia, and Strain (GRBAS). However, these metrics are subjective and prone to inter-rater variability, motivating the need for automated, objective assessment methods. This study proposes Voice Quality Assessment Network (VOQANet), a deep learning-based framework with an attention mechanism that leverages a Speech Foundation Model (SFM) to extract high-level acoustic and prosodic information from raw speech. To enhance robustness and interpretability, we also introduce VOQANet+, which integrates handcrafted acoustic features—such as jitter, shimmer, and harmonics-to-noise ratio (HNR)—with SFM embeddings into a hybrid representation. Unlike prior studies focused only on vowel-based phonation (PVQD-A subset) of the Perceptual Voice Quality Dataset (PVQD), we evaluate our models on both vowel-based and sentence-level speech (PVQD-S subset) to improve generalizability. Results show that sentence-based input outperforms vowel-based input, especially at the patient level, underscoring the value of longer utterances for capturing perceptual voice attributes. VOQANet consistently surpasses baseline methods in root mean squared error (RMSE) and Pearson correlation coefficient (PCC) across CAPE-V and GRBAS dimensions, with VOQANet+ achieving even better results. Additional experiments under noisy conditions show that VOQANet+ maintains high prediction accuracy and robustness, supporting its potential for real-world and telehealth deployment. These findings demonstrate the effectiveness of combining SFM embeddings with domain-informed acoustic features for interpretable and resilient voice quality assessment.
@@
Perceptual voice quality assessment, VOQANet, CAPE-V, GRBAS, speech foundation models, voice disorder assessment
@@ 1 Introduction
Voice disorders are common in modern society, and impaired voice quality can seriously affect an individual’s communication ability and social well-being [1], [2]. Voice disorders can be caused by a variety of diseases, including vocal fold nodules, polyps, paralysis, neurological diseases, such as Parkinson’s disease, and head and neck cancers [3]. These diseases often affect vocal characteristics, such as hoarseness, breathiness, roughness, or strain, which require perceptual examination by trained clinicians. Therefore, vocal signal analysis has become a non-invasive and accessible screening tool and is increasingly being adopted by otolaryngology and neurology clinics to help early detection and monitoring of voice-related disorders [4]. Auditory-perceptual voice quality assessment (VQA) aims to improve the diagnosis, monitoring, and treatment of voice disorders by providing an objective and standardized assessment of vocal function. It serves as a critical tool for identifying pathological voice conditions, tracking disease progression, and evaluating the effectiveness of therapeutic interventions. Traditionally, VQA relies on perceptual assessments by experienced clinicians using standardized scales, such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and the Grade, Roughness, Breathiness, Asthenia, Strain (GRBAS) [5],[6],[7]. CAPE-V has been adapted to multiple languages, including French, Turkish, European Portuguese, and Japanese, further demonstrating its clinical relevance and international applicability [8], [9],[10], [11]. It provides continuous ratings (0–100) for perceptual attributes, whereas GRBAS uses a discrete 4-point ordinal scale (0–3). For both, the larger the value, the more severe the condition. Although these perceptual ratings are widely used and provide valuable qualitative insights into voice disorders [12], they are inherently subjective and prone to inter- and intra-examiner variability. Recent studies have explored other strategies, such as crowdsourcing perceptual ratings, particularly in neurological disorders like Parkinson’s disease, to improve scalability and maintain rating validity [13]. The reliance on expert raters makes standardization difficult and increases the need for automated VQA.
@@
To address the above challenges, several machine learning (ML) and deep learning methods have been explored to automate perceptual voice assessment [14]. Traditional ML models, such as Random Forest (RF), Support Vector Machine (SVM), and k-Nearest Neighbors (KNN), have been used to predict CAPE-V values based on handcrafted acoustic features [15],[16]. Dispite providing interpretability, they often struggle with generalizability and fail to capture the complex characteristics of speech that influence perceptual voice quality ratings. To overcome these limitations, recent studies have employed deep learning and ensemble frameworks for pathological voice classification [2], demonstrating the benefits of combining multiple modalities and learned representations for more robust voice assessment. Recent advances in Speech Foundation Models (SFM), especially large-scale pre-trained models such as WavLM [17], HuBERT [18], and Whisper [19], have performed well in extracting high-level speech representations. WavLM and HuBERT are trained using self-supervised learning (SSL), while Whisper adopts a semi-supervised paradigm, leveraging large-scale audio-text pairing data collected through weak supervision. These models provide robust and transferable features, which are well-suited for downstream tasks, including VQA. Notably, WavLM incorporates a denoising pre-training objective, which may improve the model’s robustness in noisy clinical recordings; while Whisper’s multilingual training provides a wider range of phonetic coverage, although there may be a domain mismatch with disordered speech.
@@
In this study, we propose VOQANet (Voice Quality Assessment Network), a deep learning-based framework with an attention mechanism that leverages SFM embeddings of WavLM for perceptual VQA.To further enhance clinical relevance and model robustness, we further propose VOQANet+, an extended version that combines SFM embeddings with handcrafted acoustic features such as jitter, shimmer, and harmonics-to-noise ratio (HNR), which have been widely used in vocal pathology analysis for their ability to capture voice irregularities and instabilities  [20]. This combination enables VOQANet+ to benefit from both high-level learned representations and complementary low-level signal-based features. We evaluate our models on the Perceptual Voice Quality Dataset (PVQD), reporting both utterance-level and patient-level results, where the latter predictions are averaged across each speaker’s utterances. Experimental results show that VOQANet provides a strong baseline, while VOQANet+ consistently improves prediction accuracy and generalization, especially under noisy conditions.
@@
The main contributions of this study are summarized as follows: First, we propose VOQANet, a deep learning framework with an attention mechanism that systematically evaluates the effectiveness of SFM embeddings of pre-trained speech models for perceptual VQA. Second, we propose VOQANet+, an extended version of VOQANet that combines handcrafted acoustic features with SFM embeddings to improve model interpretability and performance by integrating domain-specific knowledge. Third, we conduct a comprehensive evaluation on the PVQD dataset, including both utterance-level and patient-level evaluations, aligning with clinical assessment practice. Through comprehensive evaluations on the PVQD dataset, our results demonstrate the potential of SFM-driven methods for robust, interpretable, and clinically relevant automated VQA.
@@ 2 Related Work
2.1 Automated Perceptual Voice Quality Assessment
Ensuring robustness and generalizability is critical for real-world applications of perceptual VQA. Traditional methods rely on machine learning models, such as RF, SVM, and KNN, which utilize handcrafted acoustic features like jitter, shimmer, zero crossing rate, and HNR [16]. A lightweight feature extraction method has been proposed to leverage these classical ML models for CAPE-V prediction. However, while these models provide interpretability and domain relevance, they often struggle to generalize across datasets due to the variability of speech patterns and the sensitivity of handcrafted acoustic features to recording conditions [16]. While CAPE-V provides continuous ratings and GRBAS uses a discrete scale, both are susceptible to inter-rater differences [21], [22], which further motivates the development of objective, automated assessment methods.
To overcome these limitations of classical methods, recent studies have adopted deep learning-based methods, especially leveraging SFMs such as Whisper [19] and WavLM [17], which learn rich acoustic representations from large-scale raw waveforms. Among these models, WavLM includes a denoising pre-training objective, which enables the model to learn robust representations even in the presence of background noise and acoustic variations. This feature is particularly beneficial for disordered speech, which often deviates from typical acoustic patterns. On the other hand, Whisper is trained on a large-scale multilingual and multitask corpus, which makes it highly effective in a variety of automatic speech recognition (ASR) tasks. Whisper has been explored for other applications, such as improving the performance of speaker verification (SV) tasks [23]. These findings highlight the potential of SFMs in clinical applications, while also underscoring the need to tailor the representations to domain-specific characteristics. However, SFM embeddings are typically learned from general speech corpora and may not fully capture task-specific or clinically salient characteristics relevant to voice pathology. To overcome this limitation, recent studies have explored hybrid modeling strategies that combine the generalization capabilities of SFM embeddings with handcrafted features to improve the prediction performance of VQA systems. This has shown that such hybrid audio representations can simultaneously enhance robustness and interpretability in cognitively and physically demanding speech tasks  [24]. Based on these insights, we explore deep learning-based methods that combine SFM representations with clinically relevant feature representations.
@@ 2.2 Speech Foundation Models
Models pre-trained on large-scale corpora such as Librispeech  [25] have achieved excellent performance in various speech tasks such as ASR, SV, speech synthesis, and speech emotion recognition [26], [27]. In the context of VQA, SFMs have been explored for their ability to extract rich and contextualized acoustic representations that can be related to perceptual voice attributes such as breathiness, strain, and roughness. These deep representations are able to capture complex speech patterns that are usually difficult to model using only handcrafted acoustic features.
2.3 Hybrid Models Combining SFM and Handcrafted Features
Combining SFM embeddings with handcrafted acoustic features has been explored to capture data-driven and clinically interpretable speech characteristics. A hybrid approach combining data-driven features from a BYOL-derived model with handcrafted features extracted using openSMILE has demonstrated strong performance on speech analysis tasks [28]. Similarly, self-supervised speech models jave been shown to outperform traditional acoustic features in predicting parameters such as heart activity, highlighting the efficacy of combining SSL embeddings with handcrafted features [24]. While prior studies have explored hybrid designs in affective or biomedical speech tasks, few have systematically examined such approaches in the context of clinical voice assessment, particularly with respect to GRBAS and CAPE-V prediction. Although clinical findings suggest that CAPE-V ratings may correlate more strongly with objective acoustic and aerodynamic measures than GRBAS in certain populations [29], both scales remain widely used in perceptual VQA and offer complementary insights. Therefore, we include both in this study.
@@ 2.4 Evaluation Strategies in Voice Assessment Models
Assessment strategies play a critical role in evaluating the reliability and clinical applicability of voice assessment models. Most previous studies on automated VQA focus on predicting perceptual scores at the utterance level, where each audio segment is treated as an independent sample [16, 28]. While this approach enables fine-grained analysis, it may not capture broader patterns of the patient’s overall vocal characteristics, especially when there are phonetic differences between different utterances. To improve reliability in clinical settings, some studies have begun to explore patient-level assessments by aggregating predictions from multiple utterances from the same individual [24]. This strategy better reflects the real-world diagnostic process, where clinicians assess voice quality based on a complete set of speech samples rather than isolated fragments.
@@ 3 Proposed Method
The overall architectures of VOQANet and VOQANet+ are shown in Fig. 1. VOQANet only uses SFM embeddings as input, while VOQANet+ combines SFM and handcrafted features as input.
@@ 3.1 VOQANet: SFM-Based Feature Learning
where $h^{(\ell)}$ represents the hidden state at layer $\ell$ of the pre-trained SFM, and $\alpha_{\ell}$ is a learnable scalar weight normalized by softmax. This layer-wise aggregation strategy follows prior work that showed that weighted combinations of intermediate layers outperform single-layer embeddings in speech assessment tasks [30]. This mechanism allows the model to adaptively emphasize the layers most relevant to perceptual voice quality.
@@ 3.2 VOQANet+: Joint Representation Learning
As shown in Fig. 1(b), to further enhance the model’s ability to capture clinically relevant speech characteristics, VOQANet+ combines handcrafted acoustic features with SFM embeddings. These handcrafted features include Jitter, Shimmer, and HNR, which are extracted from the same waveform $\mathbf{w}$ using signal processing techniques:
@@
where $f_{A}(\cdot)$ denotes the handcrafted feature extraction function. The final feature representation is obtained by concatenating the two feature types:
@@
As a result, a 1027-dimensional hybrid feature vector sequence is obtained. This combination enables the model to jointly learn from high-level SFM embeddings and low-level handcrafted acoustic features, thereby improving the robustness and interpretability of perceptual VQA.
@@ 4.2 Data Split for Training and Testing
To prevent data leakage and ensure generalizability, the PVQD dataset was split at the patient level, meaning that each speaker was exclusively assigned to either the training set or the test set. This ensured that the model was evaluating voice quality on unseen speakers rather than memorized speech patterns, thus validating its ability to generalize beyond the training set. This approach follows that used in [16] to ensure consistency with previous studies. The PVQD dataset originally contained 296 recordings, but 13 corrupted files were excluded, leaving a total of 283 valid samples. Specifically, 226 samples were used for training and 57 samples for testing. Table 2 provides the demographic distribution of speakers in the training and test sets, including gender ratio and age range. For a more comprehensive evaluation, vowel segments and continuous speech segments were extracted from each recording. Therefore, the PVQD dataset was divided into two subsets: PVQD-A (vowel-only subset), which contains /a/ vowel segments; and PVQD-S (speech-based subset), which consists of continuous speech segments. In this way, vowel phonation (PVQD-A) and continuous speech (PVQD-S) were evaluated independently. The number of samples in each subset is shown in Table 3. Since each recording contains one /a/ vowel and multiple continuous speech segments, there are more samples of continuous speech than /a/ vowel.

### v3

# Towards Robust Assessment of Pathological Voices via Combined Low-Level Descriptors and Foundation Model Representations

@@ Abstract
Perceptual voice quality assessment is essential for diagnosing and monitoring voice disorders by providing standardized evaluations of vocal function. Traditionally, expert raters use standard scales such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and Grade, Roughness, Breathiness, Asthenia, and Strain (GRBAS). However, these metrics are subjective and prone to inter-rater variability, motivating the need for automated, objective assessment methods. This study proposes Voice Quality Assessment Network (VOQANet), a deep learning-based framework with an attention mechanism that leverages a Speech Foundation Model (SFM) to extract high-level acoustic and prosodic information from raw speech. To enhance robustness and interpretability, we also introduce VOQANet+, which integrates low-level speech descriptors—such as jitter, shimmer, and harmonics-to-noise ratio (HNR)—with SFM embeddings into a hybrid representation. Unlike prior studies focused only on vowel-based phonation (PVQD-A subset) of the Perceptual Voice Quality Dataset (PVQD), we evaluate our models on both vowel-based and sentence-level speech (PVQD-S subset) to improve generalizability. Results show that sentence-based input outperforms vowel-based input, especially at the patient level, underscoring the value of longer utterances for capturing perceptual voice attributes. VOQANet consistently surpasses baseline methods in root mean squared error (RMSE) and Pearson correlation coefficient (PCC) across CAPE-V and GRBAS dimensions, with VOQANet+ achieving even better results. Additional experiments under noisy conditions show that VOQANet+ maintains high prediction accuracy and robustness, supporting its potential for real-world and telehealth deployment. These findings demonstrate the effectiveness of combining SFM embeddings with domain-informed acoustic features for interpretable and resilient pathological voice quality assessment.
@@
Pathological voice quality assessment, VOQANet, CAPE-V, GRBAS, speech foundation models, voice disorder assessment
@@ 1 Introduction
Voice disorders are common in modern society, and pathological voice quality can seriously affect an individual’s communication ability and social well-being [1], [2]. Voice disorders can be caused by a variety of diseases, including vocal fold nodules, polyps, paralysis, neurological diseases, such as Parkinson’s disease, and head and neck cancers [4]. These diseases often affect vocal characteristics, such as hoarseness, breathiness, roughness, or strain, which require perceptual examination by trained clinicians. Therefore, vocal signal analysis has become a non-invasive and accessible screening tool and is increasingly being adopted by otolaryngology and neurology clinics to help early detection and monitoring of voice-related disorders [5]. Auditory-perceptual assessment of pathological voice quality (VQA) aims to improve the diagnosis, monitoring, and treatment of voice disorders by providing an objective and standardized assessment of vocal function. It serves as a critical tool for identifying pathological voice conditions, tracking disease progression, and evaluating the effectiveness of therapeutic interventions. Traditionally, VQA relies on perceptual assessments by experienced clinicians using standardized scales, such as the Consensus Auditory-Perceptual Evaluation of Voice (CAPE-V) and the Grade, Roughness, Breathiness, Asthenia, Strain (GRBAS) [6],[7],[8]. CAPE-V has been adapted to multiple languages, including French, Turkish, European Portuguese, and Japanese, further demonstrating its clinical relevance and international applicability [9], [10],[11], [12]. It provides continuous ratings (0–100) for perceptual attributes, whereas GRBAS uses a discrete 4-point ordinal scale (0–3). For both, the larger the value, the more severe the condition. Although these perceptual ratings are widely used and provide valuable qualitative insights into voice disorders [13], they are inherently subjective and prone to inter- and intra-examiner variability. Recent studies have explored other strategies, such as crowdsourcing perceptual ratings, particularly in neurological disorders like Parkinson’s disease, to improve scalability and maintain rating validity [14]. The reliance on expert raters makes standardization difficult and increases the need for automated VQA.
@@
To address the above challenges, several machine learning (ML) and deep learning methods have been explored to automate perceptual voice assessment [15]. Traditional machine learning models, such as Random Forest (RF), Support Vector Machine (SVM), and k-Nearest Neighbors (KNN), have been used to predict CAPE-V scores based on low-level speech descriptors (LLDs), also known as handcrafted acoustic features [16],[17]. Dispite providing interpretability, they often struggle with generalizability and fail to capture the complex characteristics of speech that influence perceptual voice quality ratings. To overcome these limitations, recent studies have employed deep learning and ensemble frameworks for pathological voice classification [2], demonstrating the benefits of combining multiple modalities and learned representations for more robust voice assessment. To further enhance clinical VQA, attention-based models have also been proposed for predicting GRB scores from sustained vowel phonation, showing improved accuracy over earlier neural architecture [3]. Recent advances in Speech Foundation Models (SFM), especially large-scale pre-trained models such as WavLM [18], HuBERT [19], and Whisper [20], have performed well in extracting high-level speech representations. WavLM and HuBERT are trained using self-supervised learning (SSL), while Whisper adopts a semi-supervised paradigm, leveraging large-scale audio-text pairing data collected through weak supervision. These models provide robust and transferable features, which are well-suited for downstream tasks, including VQA. Notably, WavLM incorporates a denoising pre-training objective, which may improve the model’s robustness in noisy clinical recordings; while Whisper’s multilingual training provides a wider range of phonetic coverage, although there may be a domain mismatch with disordered speech.
@@
In this study, we propose VOQANet (Voice Quality Assessment Network), a deep learning-based framework with an attention mechanism that leverages SFM embeddings of WavLM for perceptual VQA.To further enhance clinical relevance and model robustness, we further propose VOQANet+, an extended version that combines SFM embeddings with LLDs, such as jitter, shimmer, and harmonics-to-noise ratio (HNR), which have been widely used in vocal pathology analysis for their ability to capture voice irregularities and instabilities  [21]. This combination enables VOQANet+ to benefit from both high-level learned representations and complementary low-level signal-based features. We evaluate our models on the Perceptual Voice Quality Dataset (PVQD), reporting both utterance-level and patient-level results, where the latter predictions are averaged across each speaker’s utterances. Experimental results show that VOQANet provides a strong baseline, while VOQANet+ consistently improves prediction accuracy and generalization, especially under noisy conditions.
@@
The main contributions of this study are summarized as follows: First, we propose VOQANet, a deep learning framework with an attention mechanism that systematically evaluates the effectiveness of SFM embeddings of pre-trained speech models for perceptual VQA. Second, we propose VOQANet+, an extended version of VOQANet that combines LLDs with SFM embeddings to improve model interpretability and performance by integrating domain-specific knowledge. Third, we conduct a comprehensive evaluation on the PVQD dataset, including both utterance-level and patient-level evaluations, aligning with clinical assessment practice. Through comprehensive evaluations on the PVQD dataset, our results demonstrate the potential of SFM-driven methods for robust, interpretable, and clinically relevant automated VQA.
@@ 2 Related Work
2.1 Automated Pathological Voice Quality Assessment
Ensuring robustness and generalizability is critical for real-world applications of perceptual VQA. Traditional methods rely on machine learning models, such as RF, SVM, and KNN, which utilize LLDs, like jitter, shimmer, zero crossing rate, and HNR [17]. A lightweight feature extraction method has been proposed to leverage these classical ML models for CAPE-V prediction. However, while these models provide interpretability and domain relevance, they often struggle to generalize across datasets due to the variability of speech patterns and the sensitivity of LLDs to recording conditions [17]. While CAPE-V provides continuous ratings and GRBAS uses a discrete scale, both are susceptible to inter-rater differences [22], [23], which further motivates the development of objective, automated assessment methods.
To overcome these limitations of classical methods, recent studies have adopted deep learning-based methods, especially leveraging SFMs such as Whisper [20] and WavLM [18], which learn rich acoustic representations from large-scale raw waveforms. Among these models, WavLM includes a denoising pre-training objective, which enables the model to learn robust representations even in the presence of background noise and acoustic variations. This feature is particularly beneficial for disordered speech, which often deviates from typical acoustic patterns. On the other hand, Whisper is trained on a large-scale multilingual and multitask corpus, which makes it highly effective in a variety of automatic speech recognition (ASR) tasks. Whisper has been explored for other applications, such as improving the performance of speaker verification (SV) tasks [24]. These findings highlight the potential of SFMs in clinical applications, while also underscoring the need to tailor the representations to domain-specific characteristics. However, SFM embeddings are typically learned from general speech corpora and may not fully capture task-specific or clinically salient characteristics relevant to voice pathology. To overcome this limitation, recent studies have explored hybrid modeling strategies that combine the generalization capabilities of SFM embeddings with LLDs to improve the prediction performance of VQA systems. This has shown that such hybrid audio representations can simultaneously enhance robustness and interpretability in cognitively and physically demanding speech tasks  [25]. Based on these insights, we explore deep learning-based methods that combine SFM representations with clinically relevant feature representations.
@@ 2.2 Speech Foundation Models
Models pre-trained on large-scale corpora such as Librispeech  [26] have achieved excellent performance in various speech tasks such as ASR, SV, speech synthesis, and speech emotion recognition [27], [28]. In the context of VQA, SFMs have been explored for their ability to extract rich and contextualized acoustic representations that can be related to perceptual voice attributes such as breathiness, strain, and roughness. These deep representations are able to capture complex speech patterns that are usually difficult to model using only LLDs.
2.3 Hybrid Models Combining SFM and LLDs
Combining SFM embeddings with LLDs has been explored to capture data-driven and clinically interpretable speech characteristics. A hybrid approach combining data-driven features from a BYOL-derived model with LLDs extracted using openSMILE has demonstrated strong performance on speech analysis tasks [29]. Similarly, self-supervised speech models jave been shown to outperform traditional acoustic features in predicting parameters such as heart activity, highlighting the efficacy of combining SSL embeddings with LLDs [25]. While prior studies have explored hybrid designs in affective or biomedical speech tasks, few have systematically examined such approaches in the context of clinical voice assessment, particularly with respect to GRBAS and CAPE-V prediction. Although clinical findings suggest that CAPE-V ratings may correlate more strongly with objective acoustic and aerodynamic measures than GRBAS in certain populations [30], both scales remain widely used in perceptual VQA and offer complementary insights. Therefore, we include both in this study.
@@ 2.4 Evaluation Strategies in Voice Assessment Models
Assessment strategies play a critical role in evaluating the reliability and clinical applicability of voice assessment models. Most previous studies on automated VQA focus on predicting perceptual scores at the utterance level, where each audio segment is treated as an independent sample [17, 29]. While this approach enables fine-grained analysis, it may not capture broader patterns of the patient’s overall vocal characteristics, especially when there are phonetic differences between different utterances. To improve reliability in clinical settings, some studies have begun to explore patient-level assessments by aggregating predictions from multiple utterances from the same individual [25]. This strategy better reflects the real-world diagnostic process, where clinicians assess pathological voice quality based on a complete set of speech samples rather than isolated fragments.
@@ 3 Proposed Method
The overall architectures of VOQANet and VOQANet+ are shown in Fig. 1. VOQANet only uses SFM embeddings as input, while VOQANet+ combines SFM and LLDs as input.
@@ 3.1 VOQANet: SFM-Based Feature Learning
where $h^{(\ell)}$ represents the hidden state at layer $\ell$ of the pre-trained SFM, and $\alpha_{\ell}$ is a learnable scalar weight normalized by softmax. This layer-wise aggregation strategy follows prior work that showed that weighted combinations of intermediate layers outperform single-layer embeddings in speech assessment tasks [31]. This mechanism allows the model to adaptively emphasize the layers most relevant to pathological voice quality.
@@ 3.2 VOQANet+: Joint Representation Learning
As shown in Fig. 1(b), to further enhance the model’s ability to capture clinically relevant speech characteristics, VOQANet+ combines LLDs with SFM embeddings. These LLDs include Jitter, Shimmer, and HNR, which are extracted from the same waveform $\mathbf{w}$ using signal processing techniques:
@@
where $f_{A}(\cdot)$ denotes the LLD feature extraction function. The final feature representation is obtained by concatenating the two feature types:
@@
As a result, a 1027-dimensional hybrid feature vector sequence is obtained. This combination enables the model to jointly learn from high-level SFM embeddings and LLDs, thereby improving the robustness and interpretability of perceptual VQA.
@@ 4.2 Data Split for Training and Testing
To prevent data leakage and ensure generalizability, the PVQD dataset was split at the patient level, meaning that each speaker was exclusively assigned to either the training set or the test set. This ensured that the model was evaluating pathological voice quality on unseen speakers rather than memorized speech patterns, thus validating its ability to generalize beyond the training set. This approach follows that used in [17] to ensure consistency with previous studies. The PVQD dataset originally contained 296 recordings, but 13 corrupted files were excluded, leaving a total of 283 valid samples. Specifically, 226 samples were used for training and 57 samples for testing. Table 2 provides the demographic distribution of speakers in the training and test sets, including gender ratio and age range. For a more comprehensive evaluation, vowel segments and continuous speech segments were extracted from each recording. Therefore, the PVQD dataset was divided into two subsets: PVQD-A (vowel-only subset), which contains /a/ vowel segments; and PVQD-S (speech-based subset), which consists of continuous speech segments. In this way, vowel phonation (PVQD-A) and continuous speech (PVQD-S) were evaluated independently. The number of samples in each subset is shown in Table 3. Since each recording contains one /a/ vowel and multiple continuous speech segments, there are more samples of continuous speech than /a/ vowel.